OpenStudy (anonymous):

What is the equation of a line that passes through the point (0, -2) and has a slope of -3? y = -3 y = -3x - 2 y = -2x - 3 y = -2

OpenStudy (anonymous):

Use point-intercept form. y-y1=m(x-x1); using point (x1,y1) and slope m

OpenStudy (anonymous):

so it would be y-(-2)=-3(x-0); now solve for y

OpenStudy (anonymous):

y+2=-3x -2 -2 y=3x-2

OpenStudy (anonymous):

y-(-2)=y+2